The weight of baby goats is believed to be Normally distributed, with a mean of 5.75 pounds. The average weight of a random sample of 20 baby goats is found to be 6.15 pounds, with a standard deviation of 0.35 pound. What is the standard error of the mean?
Answer #1

μ = 5.75

n = 20

x-bar = 6.15

s = 0.35

Standard error of the mean = s/√n = 0.35/√20 = 0.0783